OpenStudy (anonymous):

GIVE THE VERTEX and AXIS of symmetry :: f(x) = x2 - 2x - 8

OpenStudy (anonymous):

vertex \[\frac{b}{2a}=-\frac{-2}{2\times 2}=1\] is the first coordinate second is \[1^2-2\times 1-8=9\]

OpenStudy (anonymous):

thanks! .. how about the Axis? can you help w /that?

hero (hero):

x = 1

OpenStudy (anonymous):

thanks. :) i think i was confused for a sec, lol